OpenStudy (barrelracer011):

PLEASE HELP ASAP!!!!!! The equation is f(x)=65 + 80(7) Part C: Describe the steps to graph the equation obtained above on the coordinate axes. Mention the labels on the axes and the intervals.

OpenStudy (loser66):

\(f(x) = 65 +80(\color{red}{7})\)????

OpenStudy (barrelracer011):

What?

OpenStudy (loser66):

I ask you, is that 7 or x?

OpenStudy (barrelracer011):

Ok hold on the equation is f(y)=65 + 80(7) Is this better?

OpenStudy (loser66):

to me , it' s weird. If it is a correct problem, no matter x or y, 1/ open the parentheses, it will be f(x) = 625 2/ sketch it, just a horizontal line at the mark of 625 on y axis. 3/ DAT SIT

OpenStudy (barrelracer011):

ok so just make it f(x)= 62/5